OpenStudy (anonymous):

What is the inverse function of f(x) = x3 + 5?

OpenStudy (anonymous):

x3-5

sam (.sam.):

let y=x^3+5 y-5=x^3 \[x=\sqrt[3]{y-5}\] \[f^{-1}(x)=\sqrt[3]{x-5}\]

OpenStudy (anonymous):

thanx

sam (.sam.):

yw